OpenStudy (anonymous):

State whether each of the following sets have closure for + - x and division. (Natural #s), (Rational #s) (irrational #s) (whole #s)

OpenStudy (anonymous):

do you know what "closure" means?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok then what about \(\mathbb{N}\) if you add two natural numbers, do you get a natural number?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

k then it is closed under \(+\) if you subtract one natural number from another, do you always get a natural number?

OpenStudy (anonymous):

I sorta need these quick... I have 3 more pages of homework left o_O

OpenStudy (anonymous):

then lets do them quick if you subtract, do you get a natural number?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what about \(5-10\)?

OpenStudy (anonymous):

Oh... I see what u did there. lol Then it's not closed.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Multi is Closed, right

OpenStudy (anonymous):

if you multiply two natural numbers, do you get a natural number? yes, it is closed under \(\times\)

OpenStudy (anonymous):

And division can't be a negative

OpenStudy (anonymous):

division is \(2\div 7\) a natural number?

OpenStudy (anonymous):

Nope... so it's not closed

OpenStudy (anonymous):

k now since you want to be quick, rationals are closed under addition, subtraction \[\frac{a}{b}\pm\frac{c}{d}=\frac{ad\pm bc}{bd}\]

OpenStudy (anonymous):

Thank you for this!!

OpenStudy (anonymous):

also multiplication \[\frac{a}{b}\times \frac{c}{d}=\frac{ac}{bc}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

division is a bit trickier it is closed where the operation is defined, so i would say yes

OpenStudy (anonymous):

in fact the answer is "yes" closed under division

OpenStudy (anonymous):

as for irrationals, that is trickier still

OpenStudy (anonymous):

you might think they were closed under addition, but they are not for example \[2+\sqrt{3}\] is irrational , but so is \(-\sqrt3\) and if you add them you get \(2\)

OpenStudy (anonymous):

so irrationals not closed under addition, and of course therefore not closed under subtraction as for multiplication, it is clear that they are not closed take \(\sqrt{2}\times \sqrt{2}\) and get \(2\)

OpenStudy (anonymous):

likewise for division

OpenStudy (anonymous):

Thank you so much! I do have a couple other problems similar to this, they have various numbers in brackets. Not sure how to solve them. I will write a few

OpenStudy (anonymous):

as for "whole numbers" i have never in my too long life understood the distinction between "natural" and "whole" there is some argument about what is what, and i stay out of it

OpenStudy (anonymous):

go ahead and post if i know the answer i will help

OpenStudy (anonymous):

{1} {0,1} {-1,0,1} {0,2,4,6....} {1,2,3} {1,3,5} {-1,1} that's all of them, each one with the brackets is a new problem and not part of the old one. Same directions at the top, I'm supposed to State whether the following sets have closure in each 4 ways

OpenStudy (anonymous):

ok \(\{1\}\) is pretty clearly not closed under addition since \(1+1=2\) damn sometime math is not so hard

OpenStudy (anonymous):

neither under subtraction since \(1-1=0\) however, it is closed under multiplication and division right? \[1\times 1=1,\frac{1}{1}=1\]

OpenStudy (anonymous):

you got the hang of this?

OpenStudy (anonymous):

Yep, it's just I don't understand with multiple numbers in the bracket what I'm supposed to do. Plus I have a ton more homework due tomorrow morning

OpenStudy (anonymous):

\(\{-1,0,1\}\) not closed under addition right? since \(1+1=2\)

OpenStudy (anonymous):

or subtraction but if you multiply any two of those numbers you get another one also if you divide (with the understanding that you cannot divide by 0)

OpenStudy (anonymous):

Ok great so multiply and divide are closed

OpenStudy (anonymous):

even whole numbers add any two you get another even whole number subtract, no because you get \(2-2=-2\)

OpenStudy (anonymous):

multiply, yes, the product of two evens is even divide, no \(\frac{4}{4}=1\)

OpenStudy (anonymous):

{1,2,3} i don't think this is closed under anything you can check though

OpenStudy (anonymous):

likewise for {1,3,5}

OpenStudy (anonymous):

as for {-1,1} i think it is closed wrt multiplication and division

OpenStudy (anonymous):

Thank you!